OpenStudy (anonymous):

Solve the following system using the substitution method. 10x + y = –1 5x + 2y = 16

OpenStudy (anonymous):

If you put 5x=16-2y then you can substitute 10x+y=-1 with 2(5x)+y=-1 and since 5x=16-2y we get: 2(16-2y)+y=-1 32-4y+y=-1 -3y=-33 y=11 When y=11 this gives us: 10x+11=-1 10x=-12 x=-1.2